OpenStudy (anonymous):

Given the functions f(x) = 7x + 13 and g(x) = x + 2, which of the following functions represents f[g(x)] correctly? f[g(x)] = 7x + 15 f[g(x)] = 7x + 27 f[g(x)] = 7x2 + 15 f[g(x)] = 7x2 + 27

OpenStudy (anonymous):

someone help ??

OpenStudy (setsuna-yuregeshi):

is 7x2 supposed to be this: \[7x ^{2}\]? Or &x2 as in 7 times x times 2?

OpenStudy (anonymous):

its 7x^2 but only on the answer choices not the question.

OpenStudy (anonymous):

yes

OpenStudy (setsuna-yuregeshi):

So what do you think you are supposed to do here?

OpenStudy (anonymous):

im not sure, it says with the represent f[g(x)]

OpenStudy (anonymous):

\[given f(x)=7x+13 , g(x)= x+2\] now \[f[g(x)]=7(x+2)+13\] \[f[g(x)]=7x+14+13\] \[\therefore f[g(x)=7x+27\]

OpenStudy (anonymous):

THANKYOU

OpenStudy (anonymous):

cheers! First step of problem solving- Understanding Problem/question ( source: How to Sovle It - George Polya)